Q. two 1g carbon disks 1.00cm apart have opposite charges of equal magnitude such  that there is a 10-5N force between them .the ratio of excess electrons to total atoms on negatively charged disks is:

Grade:

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Rahul,

1g ......... 1/14 * 6.023*10^23 atoms

F= 9*10^9*q2/r2

1/9*10-5*10-9*10-4=q2

q=1/3 *10-9

no of electrons on each carbon is 1/(3*1.609) *1010

ratio= 14/(3*1.6*6.023) * 10-13